OpenStudy (anonymous):

5y=5-(-3y)-15 simplify........

OpenStudy (anonymous):

t-thank you :O

OpenStudy (anonymous):

5y = 5-(-3y)-15 5y = 5+3y-15 5y = -10+3y 2y = -10 y = -5

OpenStudy (anonymous):

:D)))))

OpenStudy (anonymous):

I believe that is right, if I did my math right. It's been awhile since I've had to simplify like that.

OpenStudy (anonymous):

thank you anyway!! i SUCK at math

OpenStudy (anonymous):

The only part I am not 100% sure on is the +3y part (I think it would be a positive 3y instead of negative due to the - outside of the paranthesis) You may want to get a second opinion

OpenStudy (anonymous):

i will check the back of the book